# Correlated phases of moat-band excitons in two-dimensional systems

### v1

@@ Abstract
We study two-dimensional systems of interacting excitons with a moat dispersion, for which the ground-state energy manifold presents a ring of discrete or continuously degenerate minima around a single point in momentum space. At low densities and for an idealized, perfectly degenerate moat, we show that the excitons undergo statistical transmutation and stabilize a chiral spin liquid. At higher densities, the moat dispersion favors Bose–Einstein condensation into states occupying multiple momenta, leading to inhomogeneous condensate phases and potentially supersolidity. We discuss the impact of band-structure warping present in realistic systems and argue that it generically stabilizes Bose-condensed phases over the chiral spin liquid, and analyze the superfluid response of the former which is unconventional due to the moat band. We also demonstrate that a proper renormalization of the exciton–exciton interaction is essential for describing these phases, and show that even purely repulsive interactions can favor inhomogeneous condensates. To further explore inhomogeneous condensate phases, we employ a Gross–Pitaevskii framework with a pseudopotential approximation and map out the resulting phase diagram. We show that the presence of degenerate dispersion minima can drive supersolidity already at weak coupling, in contrast to systems with a standard parabolic dispersion. Finally, we discuss our results in the context of real excitonic systems and argue that moat-band-induced supersolidity can be within experimental reach for realistic values of the model parameters.
@@ I Introduction
Here, we propose dilute 2D excitonic systems with a moat dispersion as a platform for realizing exotic phases of correlated bosonic matter. On a moat band, the low-energy manifold comes in the form of a ring-shaped valley in momentum space containing a highly degenerate continuum of ground states that can drive strong correlations. Such an exotic dispersion could appear for instance in systems where the electron or hole band displays a camelback feature, as two of us demonstrated in Ref. [39] in the context of a band-inverted topological insulator. One intriguing possiblity concerning the correlated states on moat bands is the formation of a Bose–Einstein condensate (BEC) phase of the dilute Wannier–Mott excitons when these can be considered almost ideal bosons. In this regard, the degenerate moat offers the possibility of realizing a coherent state where multiple momenta are simultaneously occupied. This leads to breaking of the translational symmetry and can result in a supersolid, a phase combining the solid-like elastic response to shear stress with the superfluid flow of the BEC [8, 3, 33, 4]. While the potential supersolidity of excitons has been investigated before in systems with conventional parabolic bands [29, 10, 9, 14], we show here that a moat band can lead to inhomogeneous condensates also in the limit of weak interactions, yields an unusual superfluid response, and leads to a competition with other many-body phases depending on the specific degree of degeneracy of the moat. In this regard, another exciting avenue first studied in Refs. [56, 54, 55, 53, 62] for the case of hard-core bosons, is that the high degeneracy of the moat causes statistical transmutation of the excitons. The latter are then described as fermions with a unit flux-quantum attachment which occupy the lowest Landau level corresponding to the emergent effective magnetic field, and essentially realize a chiral spin liquid (CSL) state.
@@
In this work, we present a unified framework to describe many-body phases of excitons on moat bands and obtain related phase diagrams that can guide future experimental realizations of the proposed systems. We highlight how an appropriate $T$-matrix renormalization of the bare exciton–exciton interaction is crucial to obtain inhomogeneous states and exemplify this in the case of a simple bilayer potential between dipolar excitons. We also review the theory behind the statistical transmutation of bosons on moat bands and compare the energies of the resulting CSL phases with those of the BECs for the simple bilayer model. We then argue that the phenomenon of band-structure warping, expected to be present in real systems, will typically tend to destabilize the CSL and favor the formation of a (possibly supersolid) BEC, and subsequently use a Gross–Pitaevskii framework to study excitonic supersolids across a larger parameter region. Our findings establish moat-band excitons as an exciting avenue for the exploration of correlated phases and add further interest into highly tunable 2D material platforms for the realization of novel interacting bosonic matter.
@@ III.1 Bose–Einstein condensates
Figure 2: Energy difference $\delta\equiv(\tilde{E}^{\mathrm{(LO)}}-\tilde{E}^{\mathrm{(FF)}})/N$ on the $\tilde{q}_{\mathrm{m}}$$\gamma$ plane for a quadratic moat (top) and a quartic one (bottom) for two different values of the small fixed parameter $\Lambda\equiv 2\rho^{2}\gamma$. In this regime, the ordering vector has the magnitude of the moat momentum and the two phases have vanishing kinetic energy, so the energies coincide with the interaction energies. The dashed black line marks the phase boundary between the homogeneous and inhomogeneous phases. As $\rho=\sqrt{\Lambda/2\gamma}$, one observes that a smaller $\rho$ favors the formation of an LO phase at fixed $\tilde{q}_{\mathrm{m}}$ and $\gamma$, as the transition line shifts to smaller $\gamma$ and smaller $\tilde{q}_{\mathrm{m}}$. The calculation takes into account the full $T$ matrix associated with the bilayer exciton–exciton potential of Eq. (8) and shows that the existence of a moat can give rise to translationally broken phases even in the case of weak interactions and dilute exciton densities.
@@ III.2 Composite fermion states
While the picture of many-body phases of bosons in first instance naturally leads one to think of Bose–Einstein condensation, another intriguing possibility arises in the case of a moat band. Namely, as studied in Refs. [56, 54, 55, 53, 40, 62], the bosons may undergo statistical transmutation into a many-body ground state of effective composite fermions. The fact that a moat band can cause this effect can intuitively be understood from the fact that such a band exhibits a macroscopic degeneracy, which is reminiscent of the Landau levels (LLs) in the quantum Hall effect (QHE). As shown explicitly below, we can view the bosons as fermions with a unit flux-quantum attachment generating an effective magnetic field. Then, at low enough densities, the emergent fermions can populate appropriate LLs and perfectly quench their mean-field kinetic energy due to the familiar quantum-Hall physics. With both the BEC phases and the composite fermion phases having vanishing kinetic energy, the correlation energy becomes the important point of comparison between both families of states. The aforementioned articles exclusively consider hard-core bosons, i.e., bosons interacting with a repulsive delta-function potential. In this scenario, it is clear that such a composite state of fermions will have the lowest energy at the mean-field level, because the interaction energy must vanish in the case of a contact interaction due to the Pauli exclusion principle. However, as we have argued, excitons generally do not exhibit such a short-ranged interaction, and thus we expect an interesting competition between the correlation energies which we explore in this section and the next one.
@@ III.3 Phase competition
We are now in the position to compare the energies of the CSL and BEC phases. The interaction energy of the CSL state for small $\rho$ behaves as $\rho^{3}$, thus as $n^{3/2}$. This can be seen by expanding Eq. (24), but it can be rigorously shown for all $l$ as explained in Appendix D. Thus, for small densities and at the mean-field level, the CSL state is more energetically favorable than any BEC. Indeed, as can be seen for instance from Eqs. (10) and (13), the interaction energy of a BEC phase scales linearly with $n$ at all densities (insofar as the dilute limit $na_{X}^{2}\ll 1$ is satisfied) as long as we assume that the $T$ matrix only has a weak dependence on $\rho$. In fact, we numerically find that the interaction energy of the CSL is always lower than that of the FF phase, even at high densities. However, the FF state always has almost vanishing kinetic energy, while the “kinetic energy” of the CSL state can only be fully quenched as long as $n<q_{\mathrm{m}}^{2}/2\pi$. This leads to a transition from the CSL to the FF phase (and by extension also to inhomogeneous BEC phases) as a function of the density. In Fig. 3 we present the phase diagram in the nndd plane for the bilayer model in an experimentally sensible range. While the details of the phase diagram will depend on the specific form of the interaction and single-particle dispersion, the qualitative behavior described above is expected for an arbitrary long-range interaction. We stress that this plot is only a comparison between the CSL and the FF phases, and we have not considered inhomogeneous BEC phases such as the LO state described in the previous section. While these could in principle replace the FF phase at the top right corner of the diagram, they will not significantly affect the region where the CSL is present.
@@
Let us briefly comment on the situation with a true contact interaction. The interaction energy of the CSL state in this case vanishes identically, and one would assume that the FF state always has a higher energy. While this is true at the mean-field level, it turns out that fluctuations can stabilize the FF state in a region of the parameter space, as explained in Refs. [53, 62]. This can be intuitively understood from the fact that the CSL state does not truly have zero kinetic energy; the latter is only perfectly quenched at the mean-field level where the statistical magnetic field has been set to a constant. By contrast, when $\hat{\mathbf{a}}$ is allowed to fluctuate, there can be a transition to the FF state even in the case of a point interaction. More precisely, for a fixed density, the CSL is stable at large coupling while the FF state becomes favorable at small coupling. As this is already the trend that our long-range calculation reveals (because in our case the coupling is inversely proportional to the interlayer distance), we do not expect fluctuations to result in significant qualitative changes to the phase diagram of Fig. 3.
@@ III.4 The role of band-structure warping
In the previous sections we have assumed an ideal moat, i.e., one where all states are perfectly degenerate, and have seen that this favors the CSL phase at low densities, with the FF phase overtaking at higher densities. From our theory so far, and in agreement with the existing literature, it would seem that larger moats increasingly favor the CSL phase, as the LLs of Fig. 3 get shifted to the right with increasing $q_{\mathrm{m}}$ and the density necessary to reach the $l=0$ level becomes higher. However, in real systems where one would hope to observe such excitonic many-body states, another effect is present which will tend to spoil the perfect degeneracy of the moat at large $q_{\mathrm{m}}$, namely band-structure warping due to the effects of the underlying crystal lattice [18].
Typically, the low-energy effective model for electrons and holes is written as a $\boldsymbol{k}\boldsymbol{\cdot}\boldsymbol{p}$ expansion up to quadratic order in $\boldsymbol{k}$. This often results in dispersions with perfect rotational symmetry. However, including higher-order terms almost always spoils this symmetry and produces effects reflecting the point-group symmetries of the underlying lattice. For instance, in Ref. [39] we considered excitons in a Bernevig–Hughes–Zhang model for a topological insulator with a band inversion. The latter results in an inverted particle–hole continuum, which in turn results in a moat dispersion for the excitons. While this is a good approximation to obtain the behavior of the free excitons, we must be careful when making assumptions about the many-body behavior of the system that rely on the perfect degeneracy of the emerging moat.
@@
The effect of warping becomes more significant at momenta away from the $\Gamma$ point, meaning it will be more important for larger moats. Consequently, for relatively large moats we expect a return from the CSL to the BEC phases as the warping becomes more pronounced. This is because the CSL phase hinges on the simultaneous occupation of all moat momenta and will be significantly affected by the appearance of inhomogeneities in the moat energy landscape. On the other hand, BEC phases only require a discrete number of occupied momenta and are thus almost unaffected by the warping as long as their discrete rotational symmetry is compatible with that of the ground-state dispersion. Furthermore, the presence of energy barriers between the discrete minima not only tends to favor condensation, but also plays the role of weakening the effects of fluctuations. As explained in Refs. [6, 53], on a perfectly degenerate moat these may destabilize the BEC due to the 1D-like divergence of the density of states, a problem that becomes less severe under the effects of warping. Thus, if warping is assumed to be significant, then the phase diagrams of Fig. 2 remain valid as the CSL state becomes energetically unfavorable due to a macroscopic number of excitons occupying the relative maxima along the moat contour.
@@ VIII Conclusion and outlook
In this article we have explored the possible phases of a dilute system of excitons with moat bands, where the low-energy degrees of freedom lie on a circle around some high-symmetry point in the Brillouin zone. In the assumption of a perfectly degenerate moat, we have reviewed how the excitons may undergo statistical transmutation to fermions with a flux attachment. At low densities they can then realize a chiral spin liquid phase in which the emergent fermions sit in a fully filled Landau level generated by the resulting effective magnetic field, while we have shown that long-range interactions eventually drive a transition to a Bose–Einstein condensate phase at higher densities. We then argued that the warping of the band structure due to the underlying lattice of real systems will tend to break the perfect degeneracy of the moat. As a result, the latter acquires a finite number of minima separated by an energy barrier which at low temperatures is high enough to lock the excitons to these points of lowest energy. Consequently, BEC phases become the most likely outcome in the event of excitonic condensation. We have shown that the $T$-matrix renormalization of the bare exciton–exciton interaction can lead to the emergence of periodic and supersolid exciton BECs and presented corresponding self-consistent calculations. Furthermore, the superfluid response of these supersolids is unusual due to the presence of the moat, leading to a strong directional anisotropy depending on the spatial symmetry of the periodic density pattern. We have further explored the phase diagram of excitonic condensates via a soft-core pseudopotential approximation. While in the case of a normal parabolic dispersion these periodic states can only exist in the regime of high densities and strong interactions, we have shown that a moat band can help stabilize them even in the regime of low densities and weak interaction strengths. Finally, we have justified that the supersolid region of the phase diagram should be experimentally accessible under the right conditions.

### v2

# Correlated phases of moat-band excitons in low-dimensional systems

@@ Abstract
We study two-dimensional systems of interacting excitons with a moat dispersion, for which the ground-state energy manifold presents a ring of discrete or continuously degenerate minima around a single point in momentum space. At low densities, the excitons undergo statistical transmutation and stabilize a chiral spin liquid. At higher densities, the moat dispersion favors Bose–Einstein condensation into states occupying multiple momenta, leading to inhomogeneous condensate phases and potentially supersolidity. We discuss the impact of band-structure warping present in realistic systems, which may lower the formation threshold of Bose–Einstein condensate phases. We analyze the superfluid response of the latter, which is unconventional due to the moat band. We also demonstrate that a proper renormalization of the exciton–exciton interaction is essential for describing these phases, and show that even purely repulsive interactions can favor inhomogeneous condensates. To further explore inhomogeneous condensate phases, we employ a Gross–Pitaevskii framework with a pseudopotential approximation and map out the resulting phase diagram. We show that the presence of degenerate dispersion minima can drive supersolidity already at weak coupling, in contrast to systems with a standard parabolic dispersion. Finally, we discuss our results in the context of real excitonic systems and argue that moat-band-induced supersolidity can be within experimental reach for realistic values of the model parameters.
@@ I Introduction
Here, we propose dilute 2D excitonic systems with a moat dispersion as a platform for realizing exotic phases of correlated bosonic matter. On a moat band, the low-energy manifold comes in the form of a ring-shaped valley in momentum space containing a highly degenerate continuum of ground states that can drive strong correlations. Such an exotic dispersion could appear for instance in systems where the electron or hole band displays a camelback feature, as two of us demonstrated in Ref. [21] in the context of a band-inverted topological insulator. One intriguing possiblity concerning the correlated states on moat bands is the formation of a Bose–Einstein condensate (BEC) phase of the dilute Wannier–Mott excitons when these can be considered almost ideal bosons. In this regard, the degenerate moat offers the possibility of realizing a coherent state where multiple momenta are simultaneously occupied. This leads to breaking of the translational symmetry and can result in a supersolid, a phase combining the solid-like elastic response to shear stress with the superfluid flow of the BEC [22, 23, 24, 25]. While the potential supersolidity of excitons has been investigated before in systems with conventional parabolic bands [26, 27, 28, 29], we show here that a moat band can lead to inhomogeneous condensates also in the limit of weak interactions, yields an unusual superfluid response, and leads to a competition with other many-body phases depending on the specific degree of degeneracy of the moat. Regarding the latter, an exciting avenue first studied in Refs. [30, 31, 32, 33, 34] is that the high degeneracy of the moat causes statistical transmutation of the excitons. The latter are then described as fermions with a unit flux-quantum attachment which occupy the lowest Landau level corresponding to the emergent effective magnetic field, which results in a chiral spin liquid (CSL) state that realizes excitonic topological order [35, 36]. Such state has recently been observed experimentally in imbalanced InAs/GaSb semiconductor bilayers at low densities [37].
@@
In this work, we present a unified framework to describe many-body phases of excitons on moat bands and obtain related phase diagrams that can guide future experimental realizations of the proposed systems. We highlight how an appropriate $T$-matrix renormalization of the bare exciton–exciton interaction is crucial to obtain inhomogeneous BEC phases and exemplify this in the case of a simple bilayer potential between dipolar excitons. We also review the theory behind the statistical transmutation of bosons on moat bands and compare the energies of the resulting CSL phases with those of the BECs for the simple bilayer model. We then argue that the phenomenon of band-structure warping, which can be significant in real systems, can induce a transition from the CSL to a (possibly supersolid) BEC, and subsequently use a Gross–Pitaevskii framework to study excitonic supersolids across a larger parameter region. Our findings establish moat-band excitons as an exciting avenue for the exploration of correlated phases and add further interest into highly tunable 2D material platforms for the realization of novel interacting bosonic matter.
@@ III.1 Bose–Einstein condensates
Figure 2: Energy difference $\delta\equiv(\tilde{E}^{\mathrm{(LO)}}-\tilde{E}^{\mathrm{(FF)}})/N$ on the $\tilde{q}_{\mathrm{m}}$$\gamma$ plane for a quadratic moat (top) and a quartic one (bottom) for two different values of the small fixed parameter $\Lambda\equiv 2\rho^{2}\gamma$. In this regime, the ordering vector has the magnitude of the moat momentum and the two phases have vanishing kinetic energy, so the energies coincide with the interaction energies. The dashed black line marks the phase boundary between the homogeneous and inhomogeneous phases. As $\rho=\sqrt{\Lambda/2\gamma}$, one observes that a smaller $\rho$ favors the formation of an LO phase at fixed $\tilde{q}_{\mathrm{m}}$ and $\gamma$, as the transition line shifts to smaller $\gamma$ and smaller $\tilde{q}_{\mathrm{m}}$. The calculation takes into account the full $T$ matrix associated with the bilayer exciton–exciton potential of Eq. (8) and shows that the existence of degenerate minima can give rise to translationally broken phases even in the case of weak interactions and dilute exciton densities.
@@ III.2 Composite fermion states
While the picture of many-body phases of bosons in first instance naturally leads one to think of Bose–Einstein condensation, another intriguing possibility arises in the case of a moat band. Namely, as studied in Refs. [30, 31, 32, 33, 58, 34], the bosons may undergo statistical transmutation into a many-body ground state of effective composite fermions. The fact that a moat band can cause this effect can intuitively be understood from the fact that such a band exhibits a macroscopic degeneracy, which is reminiscent of the Landau levels (LLs) in the quantum Hall effect (QHE). As shown explicitly below, we can view the bosons as fermions with a unit flux-quantum attachment generating an effective magnetic field. Then, at low enough densities, the emergent fermions can populate appropriate LLs and perfectly quench their mean-field kinetic energy due to the familiar quantum-Hall physics. With both the BEC phases and the composite fermion phases having vanishing kinetic energy, the correlation energy becomes the important point of comparison between both families of states. The aforementioned articles consider hard-core bosons, i.e., bosons interacting with a repulsive delta-function potential. In this scenario, it is clear that such a composite state of fermions will have the lowest energy at the mean-field level, because the interaction energy must vanish in the case of a contact interaction due to the Pauli exclusion principle. However, as we have argued, excitons generally do not exhibit such a short-ranged interaction, and thus we expect an interesting competition between the correlation energies which we explore in this section and the next one.
@@ III.3 Phase competition
We are now in the position to compare the energies of the CSL and BEC phases. The interaction energy per particle of the CSL state for small $\rho$ behaves as $\rho^{3}$, thus as $n^{3/2}$. This can be seen by expanding Eq. (24), but it can be rigorously shown for all $l$ as explained in Appendix D. Thus, for small densities and at the mean-field level, the CSL state is more energetically favorable than any BEC. Indeed, as can be seen for instance from Eqs. (10) and (13), the interaction energy of a BEC phase scales linearly with $n$ at all densities (insofar as the dilute limit $na_{X}^{2}\ll 1$ is satisfied) as long as we assume that the $T$ matrix only has a weak dependence on $\rho$. In fact, we numerically find that the interaction energy of the CSL is always lower than that of the FF phase, even at high densities. However, the FF state always has almost vanishing kinetic energy, while that of the CSL state can only be fully quenched as long as $n<q_{\mathrm{m}}^{2}/2\pi$. This leads to a transition from the CSL to the FF phase (and by extension also to inhomogeneous BEC phases) as a function of the density. In Fig. 3 we present the phase diagram in the nndd plane for the bilayer model in an experimentally sensible range. While the details of the phase diagram will depend on the specific form of the interaction and single-particle dispersion, the qualitative behavior described above is expected for an arbitrary long-range interaction. We stress that this plot is only a comparison between the CSL and the FF phases, and we have not considered inhomogeneous BEC phases such as the LO state described in the previous section. While these could in principle replace the FF phase at the top right corner of the diagram, they will not significantly affect the region where the CSL is present.
@@
Let us briefly comment on the differences between contact and long-range interactions. The interaction energy of the CSL state in the former case vanishes identically, and one would assume that the FF state always has a higher energy. While this is true at the mean-field level, it turns out that fluctuations can stabilize the FF state in a region of the parameter space, as explained in Refs. [33, 34]. This can be intuitively understood from the fact that the CSL state does not truly have zero kinetic energy; the latter is only perfectly quenched at the mean-field level where the statistical magnetic field has been set to a constant. By contrast, when $\hat{\mathbf{a}}$ is allowed to fluctuate, there can be a transition to the FF state even in the case of a point interaction. More precisely, below a certain density-dependent critical coupling strength, the FF state becomes favorable in front of the CSL. As this is already the trend that our long-range calculation reveals (because in our case the coupling is inversely proportional to the interlayer distance), we do not expect fluctuations to result in significant qualitative changes to the phase diagram of Fig. 3. Furthermore, it is important to realize that the equation of state of the CSL depends on the details of the interaction. In the hard-core case, fluctuations of the gauge field lead to a scaling $\mu\sim n^{2}(\ln n)^{2}$ of the chemical potential of the CSL with the density. Meanwhile, for our specific bilayer interaction we have found $\mu\sim n^{3/2}$, which arises from the interaction energy rather than the kinetic term and thus remains as the relevant scaling in the presence of gauge-field fluctuations. For these reasons, and given that phase competition in the case of long-range interactions already exists at the mean-field level, in this article we do not consider effects beyond the mean-field approach.
@@ III.4 The role of band-structure warping
In the calculations of previous sections we have assumed an ideal moat, i.e., one where all states are perfectly degenerate, and have seen that this favors the CSL phase at low densities, with the FF phase overtaking at higher densities. From our theory so far, and in agreement with the existing literature, it would seem that larger moats increasingly favor the CSL phase, as the LLs of Fig. 3 get shifted to the right with increasing $q_{\mathrm{m}}$ and the density necessary to reach the $l=0$ level becomes higher. However, in real systems where one would hope to observe such excitonic many-body states, another effect is present which can spoil the perfect degeneracy of the moat at large $q_{\mathrm{m}}$, namely band-structure warping due to the effects of the underlying crystal lattice [62].
Typically, the low-energy effective model for electrons and holes is written as a $\boldsymbol{k}\boldsymbol{\cdot}\boldsymbol{p}$ expansion up to quadratic order in $\boldsymbol{k}$. This often results in dispersions with perfect rotational symmetry. However, including higher-order terms almost always spoils this symmetry and produces effects reflecting the point-group symmetries of the underlying lattice. For instance, in Ref. [21] we considered excitons in a Bernevig–Hughes–Zhang model for a topological insulator with a band inversion. The latter results in an inverted particle–hole continuum, which in turn results in a moat dispersion for the excitons. While this is a good approximation to obtain the behavior of the free excitons, we must be careful when making assumptions about the many-body behavior of the system that rely on a significant degree of degeneracy of the emerging moat.
@@
Starting from a CSL on a fully degenerate moat and turning on a warping perturbation, the system will evolve as follows. For a very weak perturbation, the wave function of Eq. (23) remains a reliable variational ansatz and the CSL looks qualitatively the same. However, as the degree of warping increases and the associated energy shift becomes comparable to the many-body gap of the CSL, the bosons will undergo a significant reallocation on the manifold of thermally allowed energies. The system remains in a CSL phase, but Eq. (23) is no longer expected to describe the ground state of the system and the warping effect becomes nonperturbative from the point of view of this particular CSL state. In this case, one would have to find the new CSL ground-state wave function, which may for instance be a nematic CSL phase with anisotropic moat occupation as in Ref. [30]. As the warping increases further, the moat begins to behave more akin to a multivalley quadratic band and a transition to a BEC takes place. The warping energy scale at which this transition occurs is model dependent, and the details of how it behaves as a function of $n$ and $q_{\mathrm{m}}$ will depend on the specifics of the warping Hamiltonian. To determine the phase diagram precisely, a calculation taking into account the full energetics of the optimal CSL and the appropriate BEC is needed. However, the presence of a BEC at strong enough warping can be established on general grounds in view of the fact that its energy remains largely unaffected by the angular anisotropy as long as its discrete rotational symmetry is compatible with that of the warped moat. In other words, we expect that band-structure warping can lower the formation threshold of BEC phases in some systems.
On another note, the presence of energy barriers between the discrete minima also plays the role of weakening the effects of fluctuations in a BEC. As explained in Refs. [64, 33], on a perfectly degenerate moat fluctuations may destabilize the BEC due to the 1D-like divergence of the density of states, a problem that becomes less severe under the effects of warping. All in all, if the degree of warping is enough to favor the BEC phases, then the plots of Fig. 2 become the relevant phase diagrams in this scenario.
@@ VIII Conclusion and outlook
In this article we have explored the possible phases of a dilute system of excitons with moat bands, where the low-energy degrees of freedom lie on a circle around some high-symmetry point in the Brillouin zone. In the assumption of a perfectly degenerate moat, we have reviewed how the excitons may undergo statistical transmutation to fermions with a flux attachment. At low densities they can then realize a chiral spin liquid phase in which the emergent fermions sit in a fully filled Landau level generated by the resulting effective magnetic field, while we have shown that long-range interactions eventually drive a transition to a Bose–Einstein condensate phase at higher densities. We then argued that the warping of the band structure due to the underlying lattice of real systems will tend to break the perfect degeneracy of the moat. As a result, the latter acquires a finite number of minima separated by an energy barrier which at low temperatures is high enough to lock the excitons to these regions of lowest energy. When the effects of the warping become strong enough, a crossover from the CSL to BEC phases can take place. We have shown that the $T$-matrix renormalization of the bare exciton–exciton interaction can lead to the emergence of periodic and supersolid exciton BECs and presented corresponding self-consistent calculations. Furthermore, the superfluid response of these supersolids is unusual due to the presence of the moat, leading to a strong directional anisotropy depending on the spatial symmetry of the periodic density pattern. We have further explored the phase diagram of excitonic condensates via a soft-core pseudopotential approximation. While in the case of a normal parabolic dispersion these periodic states can only exist in the regime of high densities and strong interactions, we have shown that a moat band can help stabilize them even in the regime of low densities and weak interaction strengths. Finally, we have justified that the supersolid region of the phase diagram should be experimentally accessible under the right conditions.